OpenStudy (anonymous):

solve for x -4x+5(x-2)= -16+10 please help

OpenStudy (lynfran):

-4x+5x-10=-6 x-10=-6 ...

OpenStudy (anonymous):

ok so I need to find out what - 10 is -6?

OpenStudy (lynfran):

hint add 10 to both sides of the equation to find x..

OpenStudy (anonymous):

ok

OpenStudy (lynfran):

what do u get for x?

OpenStudy (anonymous):

-14

OpenStudy (lynfran):

no -6+10=?

OpenStudy (anonymous):

oh, -4

OpenStudy (lynfran):

no

OpenStudy (anonymous):

no thats wrong

OpenStudy (anonymous):

4

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can i ask you another question

OpenStudy (lynfran):

sure

OpenStudy (anonymous):

hold on im gonna be typing it or trying to link a file

OpenStudy (anonymous):